OpenStudy (anonymous):

First five terms of the sequence a_n=5-n^2?

OpenStudy (anonymous):

first one is \(a_1=5-1^2=4\) second one is \(a_2=5-2^2=1\) etc

OpenStudy (anonymous):

Thank you satellite73!